Given the cost per king-size sheet set above, and assuming the manufacturer has total fixed costs of $600,000 and estimates first-year sales will be 50,000 sets Answer the following questions:

Determine the price to consumers if the company desires a 50 percent mark-up on cost.


1
Expert's answer
2021-06-14T15:04:56-0400

Price per king sheet set will be

$28 + $12 + $10 + $3 + $16 + $14 + $5 + $8 + $15 + $30 + $15 = $ 156


Appraised first year contracts = 50,000 sets

Fixed expenditure = $600,000

Regular fixed expenditure per extra-large sheet set = 600,000/50,000 = $12

Total Price per extra-large sheet set = $156 + $12 = $168


Desired edge on contracts = 50 %

Reflect the contract price will be $100x

On the probability that verge is 50% of the contracts, which suggests edge = (50/100) *100x = 50x

At this point, price must be= $(100 - 50) = $60x

Similarly, budget = $168

→ 168 = 60x

→ x = 166/60 = 2.8


In this approach, contract budget = $(100*2.8) = $280appx.

10-14

At this point, the firm provides first to the supplier who at that point deals it to vendor. Dealer at supplies it to customers at a cost of $288

Because vendor earns the edge of 20%, thus, the edge of seller shall be = (20/100) *288 = $57.6

From now, the rate upon which dealer gets the sheet set from provider will be = $(288-57.6) = $230.4

Supplier retails the sheet set to the merchant for $230.4 and acquires the edge of 10%

→ edge of the distributor = $(10/100) *230.4 = $23.4


As a result, the budget at which supplier contracts the sheet set = $(230.4 - 23.4) = $207

In this approach, budget where the business presently contracts to the supplier shall be $207